OpenStudy (azionne):

12(x - 6) = 2(x - 9) (medal)

OpenStudy (alexandervonhumboldt2):

first distribute: 12*(x-6), the distributive property is a*(b-c)=ab-ac

OpenStudy (just_one_last_goodbye):

Alex you got this? ^_^

OpenStudy (alexandervonhumboldt2):

so xD

OpenStudy (alexandervonhumboldt2):

12*(x-6)=2*(x-9) 12x-72=2x-18 do you know what do to next ?

OpenStudy (azionne):

no

OpenStudy (alexandervonhumboldt2):

you need to add 72 and subtarct 2x from both sidesto both sides: 12x-72=2x-18 12x-72+72-2x=2x-2x-18+72 10x=54 x=54/10=5 2/5